Voyage update 5 from Evan Solly, Ice Pilot

Departed all the wonderful coastal scenery and moved east out into the Ross sea polynya

Wed 3rd: Departed all the wonderful coastal scenery and moved east out into the Ross sea polynya. An extremely fecund area with an amazing chlorophyll bloom underway. This has extended from Terra Nova out to at least 168° east turning the ocean into a greeny brown colour and clogging the bongo nets when deployed. The moderate weather continues and the sea temperature is still above 2°C. RMT trawl this morning followed by a mid-water trawl on unidentified layers in the water column. The dawn trace metal casts & bongo tow was then followed by another mid-water trawl close to the bottom on marks that proved to be silver fish as expected. Multi-beamed Clary Ridge but proved to be in not the position as charted. Deployed the DTIS on the charted area in 320m and again filmed silverfish in their natural habitat. 

Thu 4th: Mid-water trawl at 00:20 then moved further eastwards. The usual trace metal CTD casts and bongo tows followed by another mid-water trawl which caught the largest amount of silverfish to date (240kg). Shifting out of the chlorophyll bloom but sea temperature still at +1.0°C. Deployed the spar buoy to again test the MOB AIS transmitter and other equipment. Deployed the CTD rosette for water for the incubators then proceeded east to mooring near the 180° longitude. Zigzagging with multi-beam operating and searching for acoustic fish marks while the ship was in transit. 

Fri 5th: Zigzagging out towards the 180° line then turned northwards for mooring A1 near the Iselin Bank area. Completed a mid-water trawl to identify a fish mark that subsequently proved to be juvenile silverfish. Trace metal pump routine then CTD cast followed by a bongo transect this morning before passage north commenced. Arrived at the mooring site A1 at 21:00, retrieved and re-deployed the mooring by midnight. 

Sat 6th: Ranging in of the A1 mooring site was completed then two mid-water tows were undertaken before the 07:00 trace metal water sampling commenced. Bongo net was later deployed for a sample and then mid-water was deployed and washed clean of any fish residue. The main engine and propeller were checked for any creepage in the control units before passage was resumed northwards again. The occasional large iceberg about though becoming more frequent the further north we progress. 

Sun 7th: Started seeing a lot of large tabular bergs once north of 70° and these slowly increased as we progressed our passage north. Passed through the almost non-existent ice-bridge at 05:00 around 68.9°S. Lots more bergy bits and growlers between the large tabular bergs now. Strong westerlies and a bright sky made spotting of these voyage disrupters difficult. Crew meeting held today and the science teams mostly all resumed into day work instead of around the clock watches. Still deploying Argo buoys, wave rider buoys & drifter buoys at assigned latitudes as we cross them. Our coldest day today at -3.6°C, brrrrrrr. 

Mon 8th: First watch with the twin icelights operated from 23:30 until 04:00 this morning. Still the odd large iceberg visible on the radar but not many visual sightings due to the snow and fog. Trace metal and bongo casts made at 08:00. Run into our first major gale at noon as expected, northerlies of up to 60 knots recorded on the anemometer this afternoon and 69 knots about 18:00. Very rough sea and a large northwesterly swell. The centre of the low went over the top of Tangaroa dropping to 959Hpa, very low indeed. The ship coming out the other side into the confused swells made for a very uncomfortable 12 hours for all voyagers this evening. Very little sleep for anyone. Conditions had eased by midnight. 

Tue 9th: Steaming northwards still below 60°S due north of the CCAMLR area early tomorrow morning, weather conditions still easing from the west. Completed the final trace metal cast and bongo tow this morning. Passed the last iceberg at 19:30 at 61° 26.9’S, 174°29.3’E. Crew will change over to day work tomorrow and the ship clean-up will begin along with preparation for de-mob. 

The second to last iceberg.